### v1

# The Impact of Copyrighted Material on Large Language Models: A Norwegian Perspective

Abstract
The use of copyrighted materials in training generative language models raises critical legal and ethical questions. This paper presents a framework for and the results of empirically assessing the impact of copyrighted materials on the performance of large language models (LLMs) for Norwegian. We found that both books and newspapers contribute positively when the models are evaluated on a diverse set of Norwegian benchmarks, while fiction works possibly lead to decreased performance. Our experiments could inform the creation of a compensation scheme for authors whose works contribute to AI development.
1 Introduction
Generative language models have fundamentally reshaped the landscape of natural language processing (NLP), enabling the development of systems that can generate and interact with human language at an unprecedented level. This includes Norwegian, for which several large language models (LLMs) have been trained and published in the recent years using different architectures and licensing choices Kummervold et al. (2021); Kutuzov et al. (2021); Samuel et al. (2023); Liu et al. (2024).
However, the vast quantities of data required for training these models often include copyrighted materials, presenting novel challenges related to intellectual property rights and compensation. These challenges have led to numerous lawsuits across jurisdictions, fundamentally questioning the legitimacy of training models on copyrighted data without explicit permissions from content creators (Panettieri, 2024; Madigan, 2024; Weisenberger et al., 2024).
The first wave of lawsuits emerged shortly after the public release of advanced generative AI models. Content creators, including authors, visual artists, and musicians, began to express concerns about the unauthorized use of their work in training datasets. Multiple class-action lawsuits were filed in the United States, accusing prominent AI companies such as OpenAI and Meta Platforms of infringing on copyright laws by using copyrighted materials without obtaining explicit permissions. The authors argued that the unauthorized use of their works without any form of compensation or recognition undermines their intellectual property rights and jeopardizes their ability to earn a living from their creative endeavors. In Europe, a coalition of news publishers has taken legal action against Google and Meta Platforms, arguing that the use of journalistic content in training models without fair remuneration constitutes a breach of copyright and undermines the sustainability of high-quality journalism. Likewise, Norwegian rightsholder organizations representing publishing houses across the country, contacted the government in late 2023 expressing their concerns over the use of their material in training generative language models and demanding some sort of compensation were their contents to be used in the training of generative language models.
In this context, this paper describes a first attempt at empirically evaluating the impact of copyrighted content in the training of LLMs for Norwegian. We introduce a set of carefully curated datasets that are later used in the training of foundational, domain-tuned, and instruction-tuned models. We establish the proper training conditions to be able to compare models trained on the different datasets. A newly created benchmarking suite is used to evaluate the performance of each individual model and make the comparison meaningful. As a collaborative effort among several institutions, we hope the results of our investigations set the basis to guide policymaking and proper compensation schemes for authors and rightsholders in Norway.
2 Methodology
The methodology involves a comprehensive analysis that spans several stages. Initially, a diverse corpus of primarily Norwegian language data is assembled, incorporating both copyrighted and non-copyrighted materials, plus materials commonly found on the Internet. This corpus serves as the foundation for training various LLMs, each with different configurations and access levels to copyrighted content. By comparing the performance of these models across a range of linguistic and natural language processing tasks, such as text generation, translation, summarization, question-answering, sentiment analysis and more, we seek to quantify the specific contributions of copyrighted materials to the overall model quality.
To ensure robustness and reliability, the evaluation framework focuses on generation capabilities, natural language understanding, and linguistically-inspired metrics. Quantitative measures include traditional NLP metrics like accuracy, F1, BLEU, and ROUGE, which provide assessments of model accuracy and fluency. Linguistic analysis, on the other hand, involves assessing the coherence, language variability, and contextual relevance of the generated outputs. This dual approach allows for a nuanced understanding of how copyrighted materials impact the performance and utility of LLMs.
3 Data Collection
At the core of our research sits the need for guidelines and well-designed compensation mechanisms in Norway. Hence, we had to abide by Norwegian law (Åndsverkloven) for the definition of which works are protected under copyright in the country and what we are permitted to do with said works (which might differ from other jurisdictions). In line with provisions that allow research on language technology and data mining, and with the consent of the Norwegian rightsholder organizations, this study primarily relied on material legally deposited at the National Library of Norway. Materials not deposited were therefore omitted. As the institution has digitized almost all books and around 85% of the newspapers ever published in the country Nasjonalbiblioteket (2024), we believe this set of works is representative enough to draw valid conclusions from its study.
We collaborated with the National Library of Norway and the rightsholder organizations to gain access to protected materials for the building of our datasets. Others sources were also added via specific agreements with third party organizations such as the Norwegian Broadcasting Corporation (NRK), the TV channel TV2, and the newspaper conglomerates Amedia and Schibsted. Publicly available text collections like Wikipedia, datasets from the HPLT de Gibert et al. (2024) and CulturaX Nguyen et al. (2024) projects, and code in different programming languages from Lozhkov et al. (2024) were also added with the objective of setting up a realistic training scenario where using Internet crawled sources is commonplace (see Figure 1). This mixture of data allowed us to evaluate the impact of high-quality copyrighted material versus non-copyrighted alternatives. The models trained on the copyrighted materials will not be made publicly available for further use and only serve the purpose of this study.
3.1 Core Datasets
Figure 1: Treemap with the number of words contributed by each source in the extended dataset.
We followed the recipe of the Norwegian Colossal Corpus (NCC) by Kummervold et al. (2022), adapting and updating it with new up-to-date contents, re-OCRing some materials, enriching their metadata, and ensuring a uniform format and functionality across datasets. The preparation involved cleaning, deduplication, metadata tagging, and language balancing to maintain consistent representation of Norwegian, preventing other languages from overshadowing it. The corpus was divided into two main datasets: a base dataset excluding copyrighted materials [^1] Except for newspapers that fall under the Language Technology Use (Språkteknologiformål), as they were already included in other datasets such as NCC., and an extended dataset that included all collected texts, thus including all of base (see Table 1).
Dataset Documents Words
base 60,182,586 40,122,626,81
extended 125,285,547 82,149,281,266
Table 1: Number of documents and words in each of the core datasets. Words refer to whitespace-separated sub-strings.
We decided to include texts from other Scandinavian (Swedish, Danish, and Icelandic) and English sources to boost the performance of the resulting language models. To ensure that languages other than Norwegian, and primarily coming via Internet crawling, were balanced, we adapted the perplexity-based sampling strategy from De la Rosa et al. (2022) to maintain a high quality in the selected data. Instead of sampling a fixed number of documents, parameters for a Gaussian curve were calculated from 500,000-1M random documents per source, utilizing Wikipedia-based Kneser-Ney language models from Wenzek et al. (2019) and Conneau et al. (2020) [^2] Built with KenLM Heafield (2011).. We also modified the perplexity calculation to account for normalized text. These parameters then guided dataset sub-sampling to target ratios per language, reducing foreign language content while maintaining quality.
Moreover, we also built three custom perplexity models for specific Norwegian domains that proved too divergent from Wikipedia: books, newspapers, and government documents. These perplexity models were used to score each document in the datasets. Based on their perplexity scores, the documents were further divided into three segments corresponding to their quartile distribution. Documents with scores below the first quartile were classified as “good”, those between $Q_{1}$ and the third quartile as “medium”, and those above $Q_{3}$ were considered “bad”. To ensure robust training, the documents in each segment were randomized. It is also important to notice than in order to maintain the language distributions for foreign languages with respect to the amount of Norwegian texts, the total number of documents in foreign languages in the extended dataset is consequently higher and slightly different (due to the sampling strategies) than that of base; we keep the same ratios.
3.2 Domain Specific Subsets
Subset Documents Words
books 492,281 18,122,699,498
newspapers 46,764,024 9,001,803,515
books + newspapers 47,256,305 26,078,915,554
fiction books 117,319 5,287,109,366
nonfiction books 359,979 12,384,323,012
nonfiction books + newspapers 42,083,532 20,340,539,068
original books 392,887 13,352,261,605
original books + newspapers 47,156,911 22,354,065,120
translated books 96,258 4,695,814,506
Table 2: Number of documents and words in each copyrighted subset of the ‘extended’ dataset.
The copyrighted materials present in the extended dataset were further divided into groups. These subsets were carefully designed to test the effect of specific features of the copyrighted content. As illustrated in Table 2, we created subsets for books vs newspapers, fiction vs factual content, and original Norwegian vs translations.
3.3 Instruction-tuning Datasets
To align the model more closely with human objectives and assess whether instruction tuning with limited high-quality data can enhance the performance of our pre-trained models across various tasks, we built upon prior work and collected nearly 5,000 instructions annotated by research assistants formatted as (instruction, input, output) triplets. In this format, instruction refers to the directive provided by humans for the model, input is an optional field containing task-related information, and output denotes the desired response that follows the given instruction.
Model Initialization GPU/hours Accelerator
[…]
extended From scratch 50K AMD MI250X
base (warm) Mistral 7B v0.1 13.8K NVIDIA H100
extended (warm) Mistral 7B v0.1 55.6K AMD MI250X
[…]
base + nonfiction books base 7.5K AMD MI250X
base + nonfiction books + newspapers base 7.5K AMD MI250X
base + newspapers base 4.8K Google TPUv4
base + books base 4.8K Google TPUv4
base + books + newspapers base 4.8K Google TPUv4
base + original books + newspapers base 9.1K AMD MI250X
base + original books base 9.1K AMD MI250X
base + translated books base 9.1K AMD MI250X
[…]
extended instruct extended 14.2 NVIDIA H100
base (warm) instruct base (warm) 14.2 NVIDIA H100
extended (warm) instruct extended (warm) 14.2 NVIDIA H100
Table 3: Model training specifications, where Model represents the model identifier and the data used for training, Initialization represents the base model used for training, GPU/hours indicates the total GPU hours required for model training, and Accelerator represents the type of accelerator used.
4 Model Training
The training phase involved multiple models, each based on the Mistral architecture Jiang et al. (2023). The training was conducted in the following stages.
  1. 1.
    To measure the overall impact of copyrighted material and its effect in realistic scenarios, we conducted pre-training on the base and extended datasets, both from scratch and ‘warm-started’: using the existing weights of the pre-trained model Mistral 7B v0.1333https://huggingface.co/mistralai/Mistral-7B-v0.1. These four core models were trained on the same amount of data (64,000 steps of 4 million sub-word tokens each) using identical sets of hyperparameters.
  2. 2.
    To further isolate the effect of different ablations of the copyrighted materials, we continuously fine tuned the model trained on base from scratch for an extra 10,000 steps on each of the 9 domain specific subsets.
  3. 3.
    The core models were also fine tuned on the instruction data for 4 iterations to evaluate their performance on downstream tasks.
Overall, we trained 17 models (7 billion parameters each) using a total of 270,000 GPU-hours. Model training specifications are shown in Table 3.3. The infrastructure for training included the LUMI supercomputer, Idunn cluster, and Google TPUs through the Tensor Research Cloud program. [^4] To assess the deviation introduced by differences in training infrastructures and platforms across the participating institutions, each team trained a control model with 1.5B parameters based on the Llama 2 architecture. The training setups were identical, utilizing the base dataset. After comparing the validation loss curves from each team, we found that the curves were almost the same, with a deviation of less than $0.05$ in terms of the final convergence validation loss. Besides, we trained two tokenizers with the base and extended datasets separately, both with the same vocabulary size of $32,768$. After an initial test of the fertility [^5] Fertility expresses the fragmentation rate of a tokenizer and is #$\#tokens/\#words$ in one corpus. of the tokenizers, we found the difference between them was only $0.0013$. Therefore, we decided to use the same tokenizer trained with the base dataset for all the models.
5 Evaluation Framework
Evaluating generative LLMs is far from a solved problem, in particular for Norwegian, where there were few existing resources at the outset of our work. We compiled a set of 28 common NLP tasks, encompassing a range of different metrics to assess the performance of each of the models. These tasks can be grouped into 9 higher level skills:
  1. 1.
    Sentiment Analysis, here defined as binary polarity classification on both the sentence- and document-level, based on the existing NoReC datasets of professional reviews Velldal et al. (2018); Øvrelid et al. (2020).
  2. 2.
    Fairness & Truthfulness. Fairness refers to the absence of bias in the predictions and outputs of a model. Evaluating fairness ensures that the model does not favor or discriminate against particular groups based on attributes like race, gender, or ethnicity. This skill was evaluated using a newly-created dataset, [^6] https://huggingface.co/datasets/mimir-project/mimir-bias which covers a wide range of bias types, including race, religion, gender, geography, occupation, age etc. Truthfulness involves the accuracy and reliability of the information produced by the model, ensuring it generates factual and verifiable content. This skill was evaluated using the newly-created NorTruthfulQA dataset which adapts the English TruthfulQA dataset Lin et al. (2021) to Norwegian and assesses whether a model is truthful in selecting and generating answers to questions that involve common human misconceptions. [^7] https://huggingface.co/datasets/ltg/nortruthfulqa_mc and https://huggingface.co/datasets/ltg/nortruthfulqa_gen
  3. 3.
    Reading Comprehension, which measures the ability of a model to understand and interpret text. It involves answering questions about a given passage, summarizing content, or explaining the meaning of specific phrases or sentences. This skill estimates how well the model grasps the context and details in the text. It was evaluated using the existing extractive question-answering NorQUAD dataset Ivanova et al. (2023).
  4. 4.
    World Knowledge, which assesses the extent of factual information a language model has about the world. This includes historical events, geographical data, scientific facts, cultural knowledge, and more. The model should correctly answer questions or provide information based on real-world knowledge. This skill was evaluated using the newly-created NorOpenBookQA [^8] https://huggingface.co/datasets/ltg/noropenbookqa dataset, which adapts the English OpenBookQA dataset Mihaylov et al. (2018) to Norwegian and the NRK-Quiz-QA [^9] https://huggingface.co/datasets/ltg/nrk dataset, consisting of a collection of online quizzes on a number of different topics.
  5. 5.
    Commonsense Reasoning, which involves the ability of a model to make logical inferences based on everyday knowledge and understanding of the world. The model should reason about situations that require practical, everyday knowledge that people take for granted. This skill was evaluated using the newly-created NorCommonsenseQA dataset [^10] https://huggingface.co/datasets/ltg/norcommonsenseqa of multiple-choice commonsense question answer-pairs which adapts the corresponding English CommonsenseQA dataset Talmor et al. (2019) to Norwegian.
  6. 6.
    Norwegian Language evaluation focuses on the ability of a model to understand and generate text in Norwegian, specifically its grammar, structure, and sentence construction. This skill is important for assessing how well the model handles Norwegian and their specific syntactic rules. It was evaluated using the existing NCB Farsethås and Tjøstheim (2024) and ASK-GEC Jentoft (2023) datasets, and the newly-created NorIdiom [^11] https://huggingface.co/datasets/mimir-project/noridiom dataset.
  7. 7.
    Summarization, which measures the ability of a model to condense longer pieces of text into shorter, coherent summaries that capture the main points. This skill is crucial for applications where users need a quick understanding of large volumes of information, such as news articles or research papers. It was evaluated using the MediaFutures datase. [^12] https://huggingface.co/datasets/SamiaT/NorSumm
  8. 8.
    Translation, which assesses how accurately a language model can convert text from one language to another while preserving the meaning, tone, and context. It was evaluated using the existing Tatoeba dataset Tiedemann (2020).
  9. 9.
    Variation and Readability, which consists of measuring the lexical diversity of a model by looking at the amount of redundancy in the text it produces and at the readability of these texts measured by average sentence length and the proportion of long words. As such, this skill evaluation did not require any specific benchmarking datasets.
[A complete description of the evaluation framework is available in a forthcoming publication which is currently under review.] All the newly created datasets will be made publicly available upon paper acceptance. We follow the standard in-context learning evaluation design for pretrained decoder-only language models (e.g., Brown et al., 2020; Touvron et al., 2023), which includes zero-shot and few-shot evaluation.
In this paper, for the sake of simplicity, we selected the most common metrics per task and aggregated scores using a simple cumulative sum per higher-level skill. In order to aggregate results into an overall score, with the caveats of aggregating metrics of different nature, scores were extracted for the best available {0, 1, 4, 16}-shot configuration for each task and the best score for each of the prompts (we formulated five different prompts for each task, but the evaluation is based on the best performing prompt for each of them). Metrics were normalized to exhibit the same higher-is-better behavior in a range of 0 to 100.
base (warm)extended (warm)base + translated booksbase + original books + newspapersbase + original booksbase + nonfiction books + newspapersbase + nonfiction booksbase + fiction booksbase + books + newspapersbase + newspapersbase + booksbaseextended441.83413.98427.30440.35435.64408.20427.40440.97427.87436.81409.54482.98480.28Sentiment AnalysisFairness & TruthfulnessReading ComprehensionWorld KnowledgeCommonsense ReasoningNorwegian LanguageSummarizationTranslationVariation & Readability
Figure 2: Total summed scores across all skills averaged by task for each model. Best scores among from-scratch models underlined, best overall from-scratch in bold. Dashed line at the base score.
6 Results
The evaluation of the trained models demonstrated that incorporating copyrighted materials provided a measurable boost in performance across a range of NLP tasks. To illustrate the overall performance differences, Figure 2 shows the total scores across all skills, averaged by task for each model. Non-aggregated scores for all tasks and models are available at the Mímir repository. [^13] https://github.com/mimir-project/mimir-evaluation-suite
6.1 Core Models
As shown in Table 4 and Figure 3, the performance analysis of core models across various tasks reveals distinct strengths for different configurations. The base (warm-started) configuration consistently excels in Sentiment Analysis, World Knowledge, and Norwegian Language. In contrast, the extended (warm-started) configuration leads in Fairness & Truthfulness, Reading Comprehension, Commonsense Reasoning, Translation, and Variation & Readability, indicating its robust performance for language-intensive tasks. The base configuration generally lags behind others, scoring the lowest across multiple tasks. Meanwhile, the extended configuration performs well, particularly in Summarization. Furthermore, it indicates that we could leverage the existing metadata available at the National Library to tailor subsets of the copyrighted material and build models that excel at specific tasks. However, the difference between the base and extended warm-started models is very small. Further testing is required to assess whether this difference is still statistically significant.
baseextendedbase(warm)extended(warm)Mistral 7Bv0.1448.01482.98480.28441.83413.98Sentiment AnalysisFairness & TruthfulnessReading ComprehensionWorld KnowledgeCommonsense ReasoningNorwegian LanguageSummarizationTranslationVariation & Readability
Figure 3: Cumulative score of the core pre-trained models under different regimes. Including original Mistral 7B v0.1 for reference.
Model SA FT RC WK RC NL S T VR
extended 3 2 3 3 2 2 1 3 2
base 4 3 4 4 3 4 3 4 3
extended (warm) 2 3 1 2 1 1 2 1 1
base (warm) 1 1 2 1 1 3 2 2 4
Table 4: Results for ranking the core models on all tasks by skill via (i) finding the best k-shot configuration for each task and (ii) aggregating metric-wise rankings. SA=Sentiment Analysis. FT=Fairness & Truthfulness. RC=Reading Comprehension. NL=Norwegian Language. WK=World Knowledge. CR=Commonsense Reasoning. S=Summarization. T=Translation. VR=Variation & Readability. Lower is better.
basebase + fiction booksbase + booksbase + nonfiction booksbase + newspapersbase + nonfiction books + newspapersbase + original books + newspapersextended6.73%5.51%6.52%6.37%3.24%3.22%-1.40%
Figure 4: Average percentage gains over the performance of the base model. Negative results indicate a decrease in performance over base, positive results a gain.
While warm-started models generally outperformed models trained from scratch, there was reduced sensitivity to the presence of copyrighted materials. This suggests that the pre-existing weights, which were primarily trained on English data, diminished the impact of adding high-quality Norwegian copyrighted texts (see more on that in 7).
6.2 Domain-tuned Models
To further explore the specific effects of different types of training data, we analyzed the gains in performance by focusing on different sub-corpora, such as newspapers, books, and mixed datasets. Figure 4 provides an overview of the average percentage gains for models trained on various data configurations compared to the base model. It shows that the extended model exhibits the highest average gain at 6.73%, indicating substantial overall improvement. The addition of nonfiction books and newspapers follows with a 6.52% gain, and the addition of only newspapers shows a 6.37% improvement. Other configurations, such as adding original books and newspapers or nonfiction books, also demonstrate positive gains of 5.51% and 3.22%, respectively. Conversely, the addition of fiction books is the only one to show a negative performance, with a decrease of 1.40%. Interestingly, when decomposed by skill, the addition of fiction books makes the model excel at generating more diverse texts.
6.3 Instruction-tuned Models
Lastly, as shown in Figure 5, when the core models are further fine-tuned on data to follow instructions, the gains across models are all consistent, showing that the core advantage lies in the pre-training data, while further training on instructions gives a consistent boost in performance. Moreover, instruction tuning also seems to reduce the gap between the base and extended configurations, suggesting that copyrighted material might become less relevant as supervised finetuning datasets increase in size in the post-training phases of LLMs. Interestingly, adding Norwegian instruction data on top of the base model seems enough to improve over the performance of Mistral 7B v0.1.
7 Discussion
Our findings underline the value of copyrighted materials in improving the performance of generative language models, particularly for specialized NLP tasks in Norwegian. The inclusion of curated, high-quality texts provides a substantial advantage in terms of language richness, coherence, and context-specific understanding. At the same time, these advantages are much less pronounced for the models warm-started with the weights pre-trained on other languages (mostly English). We see two possible reasons for this:
  1. 1.
    The amount of training data matters more than its quality or licensing status. Warm-started models are effectively trained on more data than the ‘from-scratch’ models, and at some point adding even more data brings diminishing returns (with a given model size).
  2. 2.
    Copyrighted Norwegian data is indeed beneficial for LLMs, but the original models used for warm-starting were presumably already pre-trained on parts of this data. The exact composition of the training sets of models like Mistral is not known; it is quite probable that they may contain copyrighted content. Thus, continuous pre-training on (partially) the same content did not bring the expected benefits to the warm-started extended models.
extendedextended instructbasebase instructextended (warm)extended (warm) instructbase (warm)base (warm) instructMistral 7B v0.1448.01503.36480.28503.89482.98445.39413.98458.36441.83Sentiment AnalysisFairness & TruthfulnessReading ComprehensionWorld KnowledgeCommonsense ReasoningNorwegian LanguageSummarizationTranslationVariation & Readability
Figure 5: Total scores (sum) of all averaged scores per skill for the core models and their instruct versions, with original Mistral 7B v0.1 for reference. Dashed line at the base score. Best scores in bold, second best underlined.
7.1 Ethical and Legal Considerations
The use of copyrighted materials in model training raises significant ethical and legal questions. The observed improvements in model quality must be balanced against the rights of content creators, who have not consented to the use of their work. This highlights the need for clear guidelines and fair compensation mechanisms that recognize the value of copyrighted materials in LLM development.
7.2 Implications for Policy
The empirical evidence gathered in our research is crucial for informing copyright policy in the digital age. Policymakers can use these findings to establish frameworks that ensure creators are adequately compensated, balancing the needs of LLM innovation with the rights of authors and publishers. This is particularly relevant in light of ongoing lawsuits against major AI companies, underscoring the necessity of legal clarity.
8 Conclusion
Our study represents a pioneering effort to quantify the impact of copyrighted materials on LLMs for Norwegian. Our results indicate that high-quality copyrighted content significantly enhances model performance, especially for complex NLP tasks. However, these benefits bring forth ethical and legal challenges that must be addressed to ensure a sustainable and fair approach to LLM development. By providing empirical evidence, we aim to contribute to the ongoing discourse on the role of copyright in AI and inform future policies that support both innovation and the rights of content creators.
9 Future Work
Future work should focus on testing models at various scales and different pre-trained open weights to better understand how dataset composition affects performance. By experimenting with models of different sizes, we could identify any scaling thresholds where the impact of copyrighted material varies significantly.
Additionally, the observed effects of fiction on model performance highlight the need to 1) examine how different types of fiction—such as fantasy or historical fiction—impact tasks like Sentiment Analysis and Commonsense Reasoning, and 2) design new and adequate benchmarks for evaluating the contribution of fiction in Norwegian LLMs for tasks such as creative writing, plot understanding, or descriptive language use. This targeted investigation could clarify the role of fiction in model training and help refine data curation strategies.
Lastly, exploring genre-specific influences more deeply, including essays, technical writing, and narrative nonfiction, may reveal distinct benefits or biases tied to each genre. Analyzing these nuances, even in a diachronic manner, will guide balanced genre representation in datasets and support the development of better performing models.
10 Distribution
The base dataset and models were intended to be freely distributed, as all materials included were granted redistribution permissions under different agreements. After we communicated the results of our investigations to the different partners, some rightholders demanded a reinterpretation of the agreements in the light of the results and this new era of generative AI. This will most likely prohibit us from sharing publicly the exact models trained within the Mímir project, but we will publish a Norwegian language model trained on a version of the base dataset with this material excluded.